OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. 2x – 3y = -1 y = x – 1 Answer (3, 4) (2, 3) (4, 3) Inconsistent

OpenStudy (anonymous):

Since y = x - 1, substitute into the first equation, you get 2x - 3(x - 1) = -1 2x -3x + 3 = -1 -x + 3 = -1 -x = -4 x = 4 Since x = 4, substitute that into y = x - 1, to get y = 3 So x = 4 and y = 3 is the solution.

OpenStudy (anonymous):

@Easyaspi314 Thanks for your help and your explanation. ((:

OpenStudy (anonymous):

Welcome.